OpenStudy (anonymous):

Use the table. a. How many possible pairs of jeans are there if each pair has one style and one color? b. Suppose you have one pair of jeans of each possible style and color in the table. What is the probability of choosing a pair of black jeans at random? Style Color regular light blue loose fit indigo boot cut washed slim fit black blue

OpenStudy (anonymous):

A. for each style there are 5 colors so 4*5 = # of choices , B. How many black pairs of jeans can there be , divide that by answer to A .

OpenStudy (anonymous):

okay well ive been trying to figure this out and i thought it was 4 pairs 1/5 so is it .... 20 pairs 1/5?

OpenStudy (anonymous):

Yes, it is. :)

OpenStudy (anonymous):

Find the sample space for tossing 2 coins. Then find P(exactly 1 head). can you help with this

OpenStudy (anonymous):

I think I can, first we have to figure out all the possibilities TT TH HH HT (I'm 80% sure this one still counts) Then how many of them have one Heads or "H", and divide that by how many possibilities 1.No 2.Yes 3.No 4.Yes 2/4 = 1/2

OpenStudy (anonymous):

Really? I thought it was that but I second guessed myself and said 1/8

OpenStudy (anonymous):

Yep, there is only 2^c possibilities.